# RetroGFN: Diverse and Feasible Retrosynthesis using GFlowNets

### v1

@@ 3.1 Single-Step Retrosynthesis
Figure 2: The accuracy of rejecting challenging non-feasible reactions. By definition, the top-k accuracy rejects all non-feasible reactions, while round-trip accuracy treats some portion of them as feasible as it assesses the feasibility with a machine learning model. We observe that even for the most tricky non-feasible reactions, the round-trip accuracy will likely not consider them feasible. Error bars are binomial proportion confidence intervals for $p=0.05$.
3.2 Limitations of Top-k Accuracy
@@
Top-k accuracy is one of the most widely used metrics in retrosynthesis. To calculate it for the entire dataset, we first compute the support function $\mathrm{F_{ACC}}$ for every product $p$, which informs whether the ground-truth reaction was found in top $k$ results returned by the model $f$:
$\mathrm{F_{ACC}}(f,p,k)=\mathbb{1}[\exists_{i\leq k}(f(p)_{i},p)\in D],$ (1)
@@
where $f(p)_{i}$ is the $i$-th set of reactants proposed by the model for product $p$. Top-k accuracy denotes the portion of ground-truth reactions that were retrieved by the model and can be written as $\mathrm{ACC}(f,k)=\frac{1}{n}\sum_{i=1}^{n}\mathrm{F_{ACC}}(f,p_{i},k)$.
Top-k accuracy works under the assumption that all sensible reactions for a given product are contained in the dataset. However, since there are often many different ways to make product, and it would be too expensive to try all of them. Therefore, real datasets are highly incomplete. In particular, it turns out that this assumption is not true for the USPTO-50k dataset [49, 24] which is the most widely used benchmark in the retrosynthesis community. To showcase that, we gathered all reactions returned by any of the considered retrosynthesis models (see Section 5) that are not included in USPTO-50k. Among 8409 reactions ranked top-1 by any considered model, 76 of them can be found in the USPTO-MIT dataset [25]. While top-k accuracy ignores these feasible reactions, the round-trip accuracy can account for a significant portion of them (see Figure 2).
3.3 Relevance of Round-Trip Accuracy
@@
The top-k round-trip metric uses the wider notion of feasibility than top-k accuracy. For a single product, the top-k round-trip accuracy value denotes the percentage of feasible reactions among top $k$ reactions returned by the machine learning model. The feasibility is estimated with a forward reaction prediction model $F$ which is a fine-tuned Chemformer [34] (see Section 5.1). The exact formula for top-k round-trip accuracy calculated on a product $p$ and retrosynthesis model $f$ is given by:
$\mathrm{F_{Round}}(f,p,k)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{1}[p\in F(f(p)_{i})],$ (2)
@@
where $F(f(p)_{i})$ is the set of products predicted by the forward model $F$ for a set of reactants $f(p)_{i}$. In other words, the metric measures how many reactions proposed by a backward model $f$ can be back-translated by a forward model $F$. We report the top-k round-trip accuracy for the entire dataset $D$, which can be written as $Round(f,k)=\frac{1}{n}\sum_{i}^{n}\mathrm{F_{Round}}(f,p_{i},k)$. Therefore, round-trip accuracy assesses both the diversity and feasibility of the returned reactions.
@@
Figure 2 shows the ratio of feasible reactions ignored by top-k accuracy and round-trip accuracy. The metrics were computed on the USPTO-50k dataset and the "real" feasibility was assessed with USPTO-MIT. Note that the number of ignored feasible reactions is highly underestimated as USPTO-MIT is by no means exhaustive. The space of all reactions is enormous and even simple manipulations of leaving groups of reactants (e.g. changing Cl to Br) are likely to result in a lot of feasible reactions that were not screened in the wet lab before. While all the feasible reactions cannot be included in the dataset directly, the round-trip accuracy can account for some portion of them by leveraging the generalization properties of deep learning. The fact that our round-trip accuracy can account for strictly more feasible reactions than top-k accuracy is essential in the context of drug design, even at the cost of the increased number of non-feasible reactions accounted as feasible. It is because the profit lost caused by the inability to synthesize a drug is drastically higher than the cost of performing an unsuccessful synthesis experiment [50].
@@ 3.4 Reliability of Round-Trip Accuracy
To assess the reliability of the round-trip accuracy, we want to estimate what percentage of non-feasible reactions the round-trip accuracy will treat as feasible (we call this metric acceptance accuracy). The problem with constructing a set of non-feasible reactions is that they are very rarely reported in the literature. All reaction from USPTO-50k and USPTO-MIT datasets that we consider in that paper consists only of feasible reactions. Therefore, to obtain non-feasible reactions, we assume that for every set of reactants from the USPTO-MIT all of its possible outcomes were reported. Under this assumption, we can create a non-feasible reaction by taking a set of reactants from USPTO-MIT and a product that is not a possible outcome. We create an initial set of such reactions by applying random forward templates to the $m$ reactants from USPTO-MIT (test split). Then we select a subset $C$ of size $m/10$ of obtained reactions so that all reactions have distinct products and reactants. Then, for every reaction $(r,p)\in C$, we gather 9 sets of reactants from USPTO-MIT with possibly high Tanimoto similarity to $r$ and add them to $C$. As a result, for every product $p$ from $C$, we have a set of 10 corresponding reactions that are non-feasible in a non-trivial way. We additionally ensure that the sets of reactants are distinct across the reactions. If a set of reactants were shared between some reactions, then only one of those reactions would be accepted by a forward model, artificially increasing the acceptance accuracy.
@@
The acceptance accuracy of round-trip accuracy (and underlying forward model) is reported in Figure 2. We see that the forward model accurately rejects even the most challenging reactions obtained by the forward reaction template application.
@@ 3.5 Generalization of Round-Trip Accuracy
Round-trip accuracy can be generalized to be able to use a wider class of machine learning models as a reaction feasibility proxy. We propose such a generalized metric in Appendix B and evaluate the baseline models on it.
4 RetroGFN
RetroGFN is a single-step retrosynthesis model, meaning it predicts a set of molecules that could react to a given target product (see Figure 3 a)). A product is represented as an annotated graph $G=(V,T,E)$, where nodes $V=\{v_{1},v_{2},...,v_{n}\}$ correspond to the molecule’s atoms along with associated atom symbols (types) $T$, and edges $E$ are bonds. Additionally, each node and edge have an associated vector of features that will be used when embedding a molecule.
4.1 Reaction Templates and Patterns
(left side of the regular expression) and a set of reactants’ patterns (right side). The atoms of the product pattern are mapped to atoms of reactants’ patterns. Reaction templates provide a strong inductive bias to the model while limiting it to a fixed set of possible transformations. However, we extend the covered reaction space by introducing a template composition process inspired by RetroComposer [37]. In this approach, we choose the reaction center where the template is going to be applied and compose a concrete template step by step using the building blocks, called patterns.
We extract the templates from the train split of USPTO-50k, following [51]. Each template is then split into product and reactant patterns (see Figure 3 b)). We denote a set of all encountered product patterns $PPS$ and an analogous set of reactant patterns $RPS$. The patterns do not include any molecular regular expression (SMARTS) and can be represented similarly to molecules - as annotated graphs.
4.2 Generation Process
  1. 1.
    The first phase determines a reaction center: a product pattern matched the product.
  2. 2.
    The second phase gathers the reactant patterns.
  3. 3.
In the end, the obtained template is applied to the given product and results in a final set of reactants. Figure 4 shows an example of the composition process while a detailed description of each phase can be found further in the section.
Figure 4: Illustration of the template composition process in RetroGFN for an input product. In the first phase, a product pattern and its concrete match to the atoms of the product is chosen. In the second phase, reactant patterns are gathered until all mappable atoms of the product pattern (highlighted red) can be mapped to mappable atoms of the reactant pattern (highlighted blue). In the third phase, the mapping between mappable product and reactant patterns is created and the obtained template is applied resulting in the reactants.
The core component of a GFlowNet model is a forward policy $P_{F}(a|s)$ describing the probability of taking action $a$ in the state $s$. The generation process samples a sequence of states and actions $\tau=(s_{1},a_{1},...,s_{k},a_{k},t)$ called a trajectory, where $t$ is a terminal state. In RetroGFN, an initial state $s_{1}$ is an input product, the intermediate states $s_{i}$ correspond to the partially constructed template, and the terminal state $t$ stores a final template along with a result of its application to the product. We group the states into three phases and the specific definition of $P_{F}(a|s)$ depends on the phase $i$:
Score functions for all the phases share a common Graph Neural Network (GNN) encoder, denoted as $\mathrm{gnn}_{1}$ that given a product $p=(V,E,T)$, embeds its nodes’ features: $\mathrm{gnn}_{1}(p)\in\mathbb{R}^{n\times d}$, where $n$ is the number of product nodes and $d$ is the embedding size. We overload the notation and let $\mathrm{gnn}_{1}(v_{j})$ denote the embedding of a product node $v_{j}\in V$. The GNN architecture we use is similar to the one from LocalRetro: a stack of MPNN layers with a single Transformer layer [52] on top. Details can be found in the Appendix A.
First Phase. A state $s$ in the first phase is an input product $p$. The action space $A^{1}(s)$ contains all possible atom matchings of product patterns from $PPS$ to the product $p$. An action $a\in A^{1}(s)$ contains the matched product pattern $pp\in PPS$ and the matched atom indices $I=\{i_{1},...,i_{m}\}$. The value of $i_{j}$ is an index of the product atom matched with $j$-th product pattern atom. To compute the $\mathrm{score}_{1}(s,a)$, we aggregate the representation of matched product’s nodes and put them into multi-layer perceptron $\mathrm{MLP}_{1}:\mathbb{R}^{d}\rightarrow\mathbb{R}$:
Second Phase. The second phase iteratively adds reactant patterns to the composed template. At the beginning of the phase, the list of reactant patterns is empty. The second phase action $a$ is a reactant pattern $rp_{j}\in RPS$ that is going to be added to the template. The $\mathrm{score}_{2}(s,a)$ concatenates the information from the previous phase and the reactant patterns collected so far (denoted as $R$) and feeds it to $\mathrm{MLP}_{2}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{|RPS|}$ that predicts the score for all the reactant patterns in $RPS$:
Here we select the $j$th score returned by the $\mathrm{MLP}_{2}$ as it corresponds to $rp_{j}$ reactant pattern from the action. Index embedding $e=\mathrm{E}_{A}(a)$ is a function that looks up the index of the element $a$ in the set $A$ and assigns the index a learnable embedding $e\in\mathbb{R}^{d}$ (e.g. $\mathrm{E}_{PPS}$ assign a unique learnable embedding to every $pp\in PPS$). At the end of this phase, we want to be sure that every atom from the product pattern can be mapped to some atom of the reactant pattern. Originally, each pattern had some atom mapping in the template it comes from (see Figure 3).
Although those explicit mappings are inadequate in the novel-composed template, we can leverage the knowledge that an atom was originally mapped. For every pattern, we construct a set of mappable atoms that consists of the pattern’s atoms that were mapped in the original template (see Figure 5). The composed template is allowed to map only the mappable atoms. We ensure that all mappable atoms in the composed template can be mapped by properly restricting the action space $A^{2}(s)$.
where $v_{i_{j}}$ is a product node matched with the $j$-th node of the product pattern, and $v_{kl}$ is the $l$-th node of the $k$-th reactant pattern from $R$. To embed the reactant pattern nodes, we introduce a GNN $\mathrm{gnn}_{2}$ with the same architecture as $\mathrm{gnn}_{1}$. The action space $A^{3}(s)$ contains all possible atom mappings. We call an atom mapping between two nodes possible when the atom symbols of the nodes are the same and neither of the nodes was previously mapped. The third phase ends when every node from the product pattern is mapped, resulting in a template that can be applied to the reaction center chosen in the first phase. The obtained reaction forms the terminal state $t$.
4.3 Training
We trained our RetroGFN with a modified version of Trajectory Balance Objective from [53], which for a trajectory $\tau=(s_{1},a_{1},s_{2},a_{2},...,s_{k},a_{k},t)$ is given with the formula:
The main difference from the original formulation comes from the fact that our RetroGFN is conditioned [27, 54, 55, 44, 42] on the product from the initial state $s_{1}$. Therefore, for every initial state, we estimate the incoming flow separately using $F(s_{1})$ function which is essentially an index embedding $F(s)=E_{P}(s)\in\mathbb{R}$ that looks up the set of training products $P$ and returns a learnable scalar (note that we only evaluate $F(s)$ during training). As a backward policy $P_{B}(a|s)$, we use a uniform distribution over the possible actions that could lead to state $s$. The reward is an exponential reward of the form $R(x)=\mathrm{exp}(\beta f(x))$ where $f$ is a feasibility proxy. It can be a machine learning model that predicts the feasibility of a reaction or an indicator of whether the forward reaction prediction model was able to backtranslate the reaction $x$. In the main part of the paper, we evaluate the former, while experiments with the latter can be found in the Appendix B. The forward model used during training is distinct from the one used in the round-trip evaluation.
4.4 Inference
$p(t)=\sum_{\tau:t\in\tau}\prod_{(s,a)\in\tau}P_{F}(a|s).$ (3)
To increase the accuracy of the estimation, we sample $K\cdot N$ trajectories. We leave the exploration of other estimation methods for future work. The details on the architecture and hyperparameters of both training and inference can be found in the Appendix A.
5 Experiments
5.1 Setup
Datasets. We compared the considered methods on two datasets: USPTO-50k, a default choice for benchmarking retrosynthesis models, and USPTO-MIT, which we use as a generalization benchmark for models trained on USPTO-50k. We used commonly used splits for both datasets [12, 25]. We refined the USPTO-MIT to ensure there is no overlap between it and the USPTO-50k train split.
Retrosynthesis Models. We compared our RetroGFN to well-known and recent state-of-the-art models: GLN [14], MEGAN [29], MHNreact [56], LocalRetro [51], RootAligned [30], RetroKNN [57], and Chemformer [34]. We used the wrappers of the original implementations and checkpoints from the Syntheseus repository [^1] https://github.com/microsoft/syntheseus. We used the evaluation procedure from Syntheseus that queries the model for 100 reactions, removes the duplicates, and truncates the list of reactions for every product to be no larger than 50. The same output was used both to calculate standard and round-trip metrics.
Forward Model. A forward (reaction prediction) model takes a set of reactants as an input and outputs a set of possible products. As a backbone, we used a pre-trained Chemformer model from [34]. We fine-tuned two forward models: Chemformer-Eval which was used to estimate the reaction feasibility in the round-trip accuracy (see Section 3) and Chemformer-Train which guided RetroGFN during the training (see Section 4). Chemformer-Train was fine-tuned on the train split of USPTO-50k, while Chemformer-Eval used both the train and test split of USPTO-50k.
Table 1: Top-k round-trip accuracy results on USPTO-50k along with the standard deviation. The best results in every column are bolded. We observe that for $k>1$ our RetroGFN consistently outperforms other methods.
method top-1 top-3 top-5 top-10 top-20 top-50
GLN 92.2 ± 26.9 86.9 ± 23.8 83.7 ± 22.9 77.5 ± 22.6 67.3 ± 22.3 47.5 ± 20.6
MEGAN 87.0 ± 33.6 80.7 ± 28.1 76.5 ± 26.2 68.5 ± 24.8 58.2 ± 23.1 41.7 ± 21.0
MHNreact 90.8 ± 28.9 85.0 ± 24.8 80.5 ± 23.8 71.8 ± 23.1 59.9 ± 22.4 38.5 ± 19.2
LocalRetro 93.0 ± 25.6 87.6 ± 22.7 84.6 ± 21.6 79.3 ± 21.0 71.0 ± 21.2 54.2 ± 21.0
RootAligned 93.2 ± 25.2 84.6 ± 23.5 78.5 ± 23.1 67.5 ± 22.4 53.1 ± 19.4 27.0 ± 10.2
Chemformer 91.4 ± 28.0 58.2 ± 28.1 41.6 ± 24.6 24.0 ± 17.5 12.8 ± 10.8 5.3 ± 5.0
RetroKNN 92.8 ± 25.9 87.1 ± 22.8 83.8 ± 21.8 78.4 ± 20.9 69.8 ± 21.0 53.0 ± 20.1
RetroGFN 91.7 ± 27.6 88.2 ± 22.8 85.6 ± 21.5 81.1 ± 20.7 74.8 ± 20.3 63.5 ± 20.0
method mrr top-1 top-3 top-5 top-10 top-20 top-50
GLN 0.6509 52.4 74.6 81.2 88.0 91.8 93.1
MEGAN 0.6226 48.7 72.3 79.5 86.7 91.0 93.5
MHNreact 0.6356 50.6 73.1 80.1 86.4 90.3 92.6
LocalRetro 0.6565 51.5 76.5 84.3 91.0 94.9 96.7
RootAligned 0.6886 56.0 79.1 86.1 91.0 93.3 94.2
Chemformer 0.6312 55.0 70.9 73.7 75.4 75.9 76.0
RetroKNN 0.6834 55.3 77.9 85.0 91.5 91.6 96.6
RetroGFN 0.6308 49.2 73.3 81.1 88.0 92.2 95.3
5.2 Results on USPTO-50k
The top-k round-trip accuracy results for USPTO-50k dataset can be found in Table 1. We observe that for $k>1$ RetroGFN consistently outperforms all the models. The absolute and relative advantage of RetroGFN over the second-best model on top-k round-trip increases with k, indicating that the model can return a large set of diverse and feasible reactions. Note that the forward model used during the training of RetroGFN was trained on a different data split than the one used for evaluation. In Table 2, we can find standard top-k accuracy results. Our method performs competitively with state-of-the-art single-step retrosynthesis models, especially for larger values of $k$ which is arguably more important for retrosynthesis search than $k=1$.
We observe that for $k>1$, our model consistently outperforms other models, producing plenty of feasible reactions. The absolute and relative advantage of RetroGFN over the second-best model on top-k FTC increases with k: from 3.6%p and 5.9% for k=3 to 9%p and 30.8% for k=50. The good results of RetroGFN on standard metrics and its exceptional performance on FTC evidence that one can greatly improve the results on FTC without sacrificing the performance on standard metrics. Interestingly, the Pearson correlation between top-k accuracy and top-k FTC for k=1 seems relatively high (corr=0.6, $p$-value=0.12), but it becomes insignificant for $k>1$, indicating that the standard accuracy metric and FTC are complimentary (one can greatly improve upon the FTC without improving on standard metrics).
5.3 Generalization Results on USPTO-MIT
5.4 Leveraging the Forward Model
In Appendix C, we study a simple model-agnostic way of leveraging the Chemformer-Train to maximize the results of the round-trip accuracy metric. While this approach significantly improves the round-trip accuracy results, it drastically decreases the standard top-k accuracy, especially for larger values of $k$. We leave the development of other methods of incorporating the Chemformer-Train model into the training pipeline for future work.
Table 3: Top-k round-trip accuracy on USPTO-MIT along with the standard deviation. The best results in every column are bolded. We observe that for $k>3$ our RetroGFN consistently outperforms other methods.
method top-1 top-3 top-5 top-10 top-20 top-50
GLN 83.5 ± 37.1 79.3 ± 31.0 76.3 ± 29.2 70.1 ± 27.3 60.3 ± 25.6 41.9 ± 21.9
MEGAN 81.9 ± 38.5 76.3 ± 31.4 72.3 ± 29.3 64.5 ± 27.0 54.2 ± 24.7 38.3 ± 21.4
MHNreact 82.8 ± 37.8 78.1 ± 30.8 74.4 ± 28.7 66.7 ± 26.7 55.2 ± 24.9 34.9 ± 20.1
LocalRetro 83.9 ± 36.7 80.3 ± 29.7 77.6 ± 27.7 72.5 ± 25.9 64.8 ± 24.6 49.1 ± 22.6
RootAligned 84.3 ± 36.4 77.6 ± 29.8 72.0 ± 27.8 62.0 ± 25.5 48.7 ± 21.2 24.8 ± 10.8
Chemformer 81.7 ± 38.7 53.5 ± 30.4 39.1 ± 25.4 23.0 ± 17.5 12.4 ± 10.5 5.1 ± 4.6
RetroKNN 83.3 ± 37.3 78.1 ± 30.2 74.1 ± 28.2 63.9 ± 27.0 44.0 ± 23.8 19.1 ± 12.0
RetroGFN 83.0 ± 37.5 80.3 ± 30.1 78.2 ± 28.0 73.9 ± 26.1 67.9 ± 24.6 56.9 ± 22.7
Table 4: Top-k accuracy results on USPTO-MIT. The best results in every column are bolded. We observe that our RetroGFN achieves competitive results.
method mrr top-1 top-3 top-5 top-10 top-20 top-50
GLN 0.4625 37.2 52.9 57.7 62.7 65.2 66.3
MEGAN 0.4647 37.0 53.4 58.8 63.7 66.8 68.9
MHNreact 0.4610 37.2 52.7 57.6 62.1 64.7 66.2
LocalRetro 0.4720 36.7 55.1 60.8 66.0 68.7 70.5
RootAligned 0.4960 40.2 56.8 61.7 66.1 68.5 69.5
Chemformer 0.4521 39.5 50.5 52.6 53.8 54.1 54.2
RetroKNN 0.4572 35.6 53.3 59.1 64.3 66.2 66.4
RetroGFN 0.4590 35.4 53.5 59.3 64.7 67.8 70.0

### v2

@@ 5.1 Single-Step Retrosynthesis
Figure 7: The accuracy of rejecting challenging non-feasible reactions. By definition, the top-k accuracy rejects all non-feasible reactions, while round-trip accuracy treats some portion of them as feasible, as it assesses the feasibility with a machine learning model. We observe that even for the most tricky non-feasible reactions, the round-trip accuracy will likely not consider them feasible. Error bars are binomial proportion confidence intervals for $p=0.05$.
5.2 Definition of Top-k Accuracy
@@
Top-k accuracy is one of the most widely used metrics in retrosynthesis. To calculate it for the entire dataset, we first compute the support function $\mathrm{F_{ACC}}$ for every product $p$, which informs whether the ground-truth reaction was found in the top $k$ results returned by the model $g$:
$\mathrm{F_{ACC}}(g,p,k)=\mathbb{1}[\exists_{i\leq k}(g(p)_{i},p)\in D],$ (2)
@@
where $g(p)_{i}$ is the $i$-th set of reactants proposed by the model for product $p$. Top-k accuracy denotes the portion of ground-truth reactions that were retrieved by the model and can be written as $\mathrm{ACC}(g,k)=\frac{1}{n}\sum_{i=1}^{n}\mathrm{F_{ACC}}(g,p_{i},k)$.
5.3 Definition of Round-Trip Accuracy
@@
The top-k round-trip metric uses the wider notion of feasibility than top-k accuracy. For a single product, the top-k round-trip accuracy value denotes the percentage of feasible reactions among the top $k$ reactions returned by the machine learning model. The feasibility is estimated with a forward reaction prediction model $f$, which Chemformer [irwin2022chemformer] fine-tuned on USPTO-50k train and test sets (see Section 4.1). The exact formula for top-k round-trip accuracy calculated on a product $p$ and retrosynthesis model $g$ is given by:
$\mathrm{F_{Round}}(g,p,k)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{1}[p\in F(g(p)_{i})],$ (3)
@@
where $f(g(p)_{i})$ is the set of products predicted by the forward model $f$ for a set of reactants $g(p)_{i}$ (we use beam size $=1$). In other words, the metric measures how many reactions proposed by a backward model $g$ can be back-translated by a forward model $f$. We report the top-k round-trip accuracy for the entire dataset $D$, which can be written as $Round(g,k)=\frac{1}{n}\sum_{i}^{n}\mathrm{F_{Round}}(g,p_{i},k)$. Therefore, round-trip accuracy assesses both the diversity and feasibility of the returned reactions.
5.4 Top-k Accuracy vs Round-Trip Accuracy
@@
Top-k accuracy works under the assumption that all sensible reactions for a given product are contained in the dataset. However, since there are often many different ways to make a product, it would be too expensive to try all of them. Therefore, real datasets are highly incomplete. In particular, it turns out that this assumption is not true for the USPTO-50k dataset [lowe2012extraction, schneider2016s], which is the most widely used benchmark in the retrosynthesis community. To showcase that, we gathered all reactions returned by any of the considered retrosynthesis models (see Section 4) that are not included in USPTO-50k. Among 8409 reactions ranked top-1 by any considered model, 76 of them can be found in the USPTO-MIT dataset [jin2017predicting]. While top-k accuracy ignores these feasible reactions, the round-trip accuracy can account for a significant portion of them (see Figure 7). The metrics were computed on the USPTO-50k dataset and the "real" feasibility was assessed with USPTO-MIT. Note that the number of ignored feasible reactions is highly underestimated as USPTO-MIT is by no means exhaustive. The space of all reactions is enormous, and even simple manipulations of leaving groups of reactants (e.g., changing Cl to Br) are likely to result in a lot of feasible reactions that were not screened in the wet lab before. While all the feasible reactions cannot be included in the dataset directly, the round-trip accuracy can account for some portion of them by leveraging the generalization properties of deep learning. The fact that our round-trip accuracy can account for strictly more feasible reactions than top-k accuracy is essential in the context of drug design, even at the cost of the increased number of non-feasible reactions accounted as feasible. It is because the profit lost caused by the inability to synthesize a drug is drastically higher than the cost of performing an unsuccessful synthesis experiment [burt2017burden]. In 5.6, we introduce a simplified drug-design pipeline and show that in 2 out of 3 scenarios optimizing round-trip accuracy leads to higher expected profits than optimizing top-k accuracy.
@@ 5.5 Reliability of Round-Trip Accuracy
To assess the reliability of the round-trip accuracy, we want to estimate what percentage of non-feasible reactions the round-trip accuracy will treat as feasible (we call this metric acceptance accuracy). The problem with constructing a set of non-feasible reactions is that they are very rarely reported in the literature. All reaction from USPTO-50k and USPTO-MIT datasets that we consider in that paper consists only of feasible reactions. Therefore, to obtain non-feasible reactions, we assume that for every set of reactants from the USPTO-MIT all of its possible outcomes were reported. Under this assumption, we can create a non-feasible reaction by taking a set of reactants from USPTO-MIT and a product that is not a possible outcome. We create an initial set of such reactions by applying random forward templates to the $m$ reactants from USPTO-MIT (test split). Then we select a subset $C$ of size $m/10$ of the obtained reactions so that all reactions have distinct products and reactants. We call $C$ a set of the most challenging non-feasible reactions. Then, for every reaction $(r,p)\in C$, we gather 9 sets of reactants $r_{i}$ from USPTO-MIT with possibly high Tanimoto similarity to $r$ and add $(r_{i},p)$ reactions to $C$. As a result, for every product $p$ from $C$, we have a set of 10 corresponding reactions that are non-feasible in a non-trivial way. We additionally ensure that the sets of reactants are distinct across the reactions. If a set of reactants were shared between some reactions, then only one of those reactions would be accepted by a forward model, artificially increasing the acceptance accuracy.
@@
The acceptance accuracy of round-trip accuracy (and underlying forward model) is reported in Figure 7. We see that the forward model accurately rejects even the most challenging reactions $C$ obtained by the forward reaction template application.
5.6 Round-Trip Accuracy in Drug Design Scenario
In the case of drug discovery, false-negative error rate $\beta_{E}$ is widely believed to have a much higher influence on the expected investment return than false-positive error rate $\alpha_{E}$ [burt2017burden]. It comes from the fact that $\beta_{E}$ is roughly proportional to the chances of missing a drug from the experimental pipeline, while $\alpha_{E}$ is roughly proportional to the chances of including a non-drug in the pipeline. And because the costs of an experiment are orders of magnitude lower than the expected profits from developing a drug, it’s very reasonable to decrease the chances of missing a drug (decrease $\beta_{E}$) at the expense of an increased number of failed experiments (increased $\alpha_{E}$).
In Section 5.6.1, we analyze the influence of $\alpha_{E}$ and $\beta_{E}$ of a retrosynthesis model on the expected income of a drug design pipeline. Further, in Section 5.6.2, we estimate the $\alpha_{E}$ and $\beta_{E}$ of models trained to optimize the top-k accuracy (ACC) and top-k round-trip accuracy (RT) metrics. In Section 5.6.3, we conclude that in 2 out of 3 scenarios, optimizing RT leads to better outcomes than optimizing ACC.
5.6.1 Drug Design Scenario
In this section, we describe a relationship between $\alpha_{E}$ and $\beta_{E}$ of the single-step retrosynthesis model and the expected investment return of a simplistic drug design pipeline.
Let’s assume that we use a single-step retrosynthesis model with error rates $\alpha_{E}$ and $\beta_{E}$ in the multi-step retrosynthesis planning pipeline (RETRO). For the sake of simplicity, let’s assume that for every molecule there is only one ground-truth synthetic route of length $n=5$. The probability that the ground-truth route can be retrieved by RETRO is roughly equal to $p_{1}=(1-\beta_{E})^{n}$ (at every synthetic step, we need to be able to recover the ground-truth reaction). If the ground-truth route can be found, RETRO outputs it in $m$ trials with a probability $p_{2}=(1-(1-\alpha_{E})^{n})^{m}$. The overall probability of retrieving a ground-truth route by RETRO in $m$ trials $p(\alpha_{E},\beta_{E},m)=p_{1}(1-p_{2})$.
Let’s further assume that RETRO proposes synthesis routes for a pool of drug candidates. The synthesis routes are then evaluated in the laboratory and successfully synthesized candidates are pushed further in the drug design pipeline, possibly becoming highly lucrative products. The expected income for $\alpha_{E}$ and $\beta_{E}$ is calculated as: $max_{m}(P\cdot p(\alpha_{E},\beta_{E},m)-C\cdot m)$. Its value depends mostly on the profit-to-cost ratio $\frac{P}{C}$, which relates the cost of running a single synthesis experiment ($C$) with an expected profit from being able to synthesize a given molecule ($P=P_{D}\cdot R_{D}$; where $P_{D}$ is the profit from developing a drug and the $R_{D}$ is the ratio of the drugs in the candidates pool). The concrete values of $P$ and $C$ are hard to estimate globally. For this reason, we evaluate three scenarios with $\frac{P}{C}$$\in\{10,100,1000\}$. The middle value was calculated based on the estimation of drug profit $P_{D}=2e9$ from [burt2017burden], drug ratio $R_{D}=1e-5$ from [sun202290] (there are $41e$ 4 compounds in the screening phase and only one of them will be a drug with a 10% chance), and synthesis cost $C=2e2$ from [koziarski2024rgfn].
We plot the expected income as a function of $\alpha_{E}$ and $\beta_{E}$ in Figure 8 for those three scenarios. Figure 8 recalls the common belief that $\beta_{E}$ influences the expected income much more than $\alpha_{E}$.
Figure 8: The estimated income of drug design pipeline as a function of false positive ($\beta_{E}$) and false negative ratio ($\alpha_{E}$) of a single-step retrosynthesis model. For scenarios with moderate and high profit-to-cost ratio (P/C), the $\beta_{E}$ is significantly more important than $\alpha_{E}$.
5.6.2 Estimates of Error Rates of ACC and RT
To assess $\beta$ of ACC, we gathered all reactions from USPTO-MIT that: 1) share a product with some ground-truth reaction from USPTO-50k, and 2) are not in USPTO-50k. There are 2532 such reactions, suggesting $\beta_{E}\approx 5\%$ for ACC. On the other hand, the forward model underlying RT was able to account for 1294/2532 of them, suggesting $\beta_{E}\approx 2.5\%$. We assume that $\alpha_{E}=0$ for ACC. Figure 7 suggests that $\alpha_{E}\approx 10\%$ for RT as the underlying forward model rejects  90% of the most tricky negative reactions. This number also reflects the top-1 forward prediction accuracy declared in the Chemformer paper [irwin2022chemformer].
5.6.3 Results for RT- and ACC-optimal models
We report the expected income for the ACC-optimal model (an idealized model that obtains perfect ACC; by definition it has $\alpha_{E}=0\%$, $\beta_{E}=5\%$) and RT-optimal model ($\alpha_{E}=10\%$, $\beta_{E}=2.5\%$) for scenarios with different profit-to-ratio (P/C) magnitudes in Table 6. We observe that in 2 out of 3 scenarios, optimizing RT leads to better outcomes than optimizing ACC.
Table 6: The expected return of investment in drug design pipeline for different profit-to-cost ratio P/C (columns) and single-step synthesis models optimized with respect to top-k accuracy (ACC) and round-trip accuracy (RT). We observe that in 2 out of 3 scenarios, optimizing RT leads to better outcomes than optimizing ACC.
model 10 100 1000
RT-optimal model 1.1e3 1.6e4 1.7e5
ACC-optimal model 1.3e3 1.5e4 1.5e5
@@ 5.7 Generalization of Round-Trip Accuracy
Standard round-trip accuracy formulation counts a reaction as feasible if the forward reaction prediction model outputs the product from the reaction given its reactants as input. Then, the feasibility is a binary feature of the reaction. In B, we propose a generalized version of round-trip accuracy that can deal with the probabilistic notion of feasibility, allowing the usage of a wider class of feasibility prediction models. We show that our RetroGFN can greatly improve this metric.